OpenStudy (anonymous):

∪−2.7<−1.9 SOLVE FOR U

Vocaloid (vocaloid):

add 2.7 to each side

OpenStudy (anonymous):

5.4 and 4.6

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

nope ∪−2.7 + 2.7 = ?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

nope.... remember the U

OpenStudy (anonymous):

Im still trying to figure it out

Vocaloid (vocaloid):

U - 2.7 + 2.7 = ?

OpenStudy (anonymous):

u=0

Vocaloid (vocaloid):

hint: A - B + B = A U - 2.7 + 2.7 = ?

OpenStudy (anonymous):

U

Vocaloid (vocaloid):

right -1.9 + 2.7 = ?

OpenStudy (anonymous):

0.8

Vocaloid (vocaloid):

right, so U < 0.8

OpenStudy (anonymous):

thank you